alculating Total Cash Flows Given the information for Marias Tennis Shop, Inc., in the previous two problems, suppose you also know that the firms net capital spending for 2017 was $640,000, and that the firm reduced its net working capital investment by $65,000. What was the firms 2017 operating cash flow, or OCF?

8. Cash Flow to Creditors The 2016 balance sheet of Marias Tennis Shop, Inc., showed long-term debt balance sheet showed long-term debt of $2.53 million. The 2017 income of $2.4 million, and the 2017 statement showed an interest expense of $187,000. What was the firms cash flow to creditors during 20172 Cash Flow to Stockholders the common stock account and $5.6 million in the additional paid-in surplus account. The 2017 balance sheet showed $595,000 and $6.18 million in the same two accounts, respectively. If the company paid out $270,000 in cash dividends during 2017, what was the cash flow to stockholders for the year? 9. The 2016 balance sheet of Marias Tennis Shop, Inc., showed $540,000 in in the nrevious two

please show all work
Answer #1

Maria's tennis shop inc

1. Operating cash flow (OCF) / Cash Flow From operating cash flow= Assets + Net capital spending + Changes in net working capital.

Operating cash flow (OCF) / Cash Flow From Assets = 0 + 640000 + 65000 = 705,000 $ (Since no information on assets is given,they are assumed to be zero).

2. cash flow to creditors = Interest paid - ( Ending Long Term Debt - Beginning Long Term Debt)

cash flow to creditors = 187,000 - (2530000 - 2400000) = 57000 $

3. Cash flow to stock holders = Dividends paid - (Ending common stock - beginning common stock) -(ending capital surplus - beginning capital surplus) + Ending treasury stock - beginning treasury stock)

Cash flow to stock holders = 270,000 - (595,000-540,000) - (6180000 - 5600000) + 0 = - 365,000 $

{ If a company raises more money from selling stock than it pays in dividends in a particular period, its cash flow to stockholders will be negative. If a company borrows more than it pays in interest and principal, its cash flow to creditors will be negative,else Positive. }